OpenStudy (anonymous):

The mass of Mars is about 6.421x10^20 metric tons. The mass of the sun is about 1.998x10^27 metric tons. About how many times the mass of Mars is the mass of the sun? Give your answer in scientific notation.

OpenStudy (anonymous):

It's asking, 'How many times would the mass of Mars 'fit into' the mass of the Sun'? Well, the mass of Mars would go into the mass of the Sun,\[\frac{m_{sun}}{m_{mars}}=\frac{1.998 \times 10^{27}}{6.421 \times 10^{20}}=3111664.85=3.112 \times 10^{6}\]in scientific notation to four significant places.

OpenStudy (anonymous):

thank you lokisan

OpenStudy (anonymous):

np